OpenStudy (anonymous):

Solve –2x^2 +3x – 9 = 0 using the quadratic formula plz

OpenStudy (anonymous):

-3+-√(3)^2-4(-2)(-9)/2(-2) -3+-√9 -72/-4 -3+-√63/4

OpenStudy (anonymous):

should i show you the likely answers this is wrong

OpenStudy (anonymous):

\[\frac{3\pm \sqrt{(3)^2-4(-2)(-9)}}{2(-2)}\] So,\[\frac{3\pm \sqrt{-63}}{-4}\]No solutions.

OpenStudy (anonymous):

Actually I should say no real roots